OpenStudy (anonymous):

A population of bacteria is introduced into a culture. the number of bacteria P can be modeled by P=500(1+4t/(50+t^2 )) where t is time in hours. Find the rate of change in population when t=2

OpenStudy (anonymous):

i assume this means take the derivative and evaluate at 2 right?

OpenStudy (anonymous):

i surmise that you are right

OpenStudy (anonymous):

ick. quotient rule time

OpenStudy (anonymous):

i did that but got some funky answer that was negative

OpenStudy (anonymous):

okay, maybe the derivative is off

OpenStudy (anonymous):

by the way the 1+4t isnt actually on top of the fraction. just the 4t is. sorry

OpenStudy (anonymous):

just 4t on top right?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

P(t)=500(1+4t/(50+t^2 )) P'(t) = 500 [(50+t^2).4 - 4t.2t]/(50+t^2)^2 by the quotient rule = 500 (-4t^2 + 200)/(t^2 + 50)^2 Hence P'(2) = 500 . (-16 + 200)/54^2 ~= 31.6

OpenStudy (anonymous):

derivative is \[ \frac{-2000 (t^2-50))}{(t^2+50)^2}\]

OpenStudy (anonymous):

what jamesj said.

OpenStudy (anonymous):

you need the steps or clear?

OpenStudy (anonymous):

i did my quotient rule backwards

OpenStudy (anonymous):

yeah i think i got it now

OpenStudy (anonymous):

lol common mistake goes away with practice

OpenStudy (anonymous):

i used to do it all the time in my other class, i took the same class in highschool, but i guess ive not had enough practice but it's the first time ive done it this whole assignment. :)